what should be tension applied to a wire of length 1m and mass 10gm, if it has to vibrate with fundamental frequency of 50 Hz
2 minutes ago
 

Grade:12

2 Answers

Arun
6 years ago
Dear Anand
mass m = 10 gm = 0.01
n = (1/2L) \sqrt(T/m)
Now
50 = 1/2 \sqrt(T/0.01)
100² = T/0.01
10000*0.01 = T
T = 100 N
Regards
Arun (askIITians forum expert)
Khimraj
6 years ago
Let tension applied to wire is T.
For fundamental frequency λ = 2L = 2*1 = 2m
u = m/L = (10*10-3)/1 = 10-2 kg/m
velocity v = (T/u)1/2
frequency f = v/λ
50 = v/2
So v = 100
(T/u)1/2 = 100
T/u = 10000
T/10-2  = 10000
So T = 100 N.
Hope it clears.